OpenStudy (anonymous):

A 9-inch tall candle burns at a rate of 0.4 inches per hour. The height of the candle at any hour is modeled by the equation: c = -0.4t + 9, where c = the height of the candle and t = the time in hours. Find the height of the candle after 2 hours. Also, how long will it take the candle to burn out?

OpenStudy (anonymous):

Height of the candle after 2 hours: where you see a "t" in the equation, replace it with the number "2". So c = -0.4 * (2) + 9 we see that (-0.4)*(2) = -0.8 so c = 9 -0.8 c = 8.2 inches. How long take for candle to burn out? This is another way of asking what is "t" when c=0. So we plug in: 0 = (-0.4)(t) +9 I'll leave it to you to solve for "t".